Circular No. 7344

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 XS_{87} (LINEAR) An object that was assumed to be asteroidal was found by LINEAR on

1999 Dec. 7 and 8 (discovery observation given below; see also MPS 8038), and it was later linked to observations by LINEAR on 2000 Jan. 6 and 7 by G. V. Williams, Minor Planet Center. Following a request from the Minor Planet Center after seeing that the orbit appeared cometlike, M. Tichý and Z. Moravec obtained observations at Kleť on Jan. 11 and 12 that showed this object to be diffuse with a coma diameter of 15′′.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 7.37830  6^{h}40^{m}43.^{s}95  +21^{o}14′14′′.5  17.3

Full astrometry and orbital elements appear on MPEC 2000-A43.

T = 1999 Aug. 6.5518 TT  ω = 151.^{o}3264
e = 0.840979  Ω = 266.7267 }2000.0
q = 2.771006 AU  i = 14.8424

a = 17.425430 AU n^{o} = 0.0135497 P = 72.740 years

SUPERNOVA 1999et IN NGC 1643 H. Yamaoka, Kyushu University; and M. Yamamoto, Ayabe Astronom-

ical Observatory (AAO), report a precise position for SN 1999et (cf. IAUC 7304), from images taken with the 0.95-m AAO reflector: α = 4^{h}43^{m}43^{s}.53, δ = −5^{o}19′19′′.8 (equinox 2000.0). During Nov. 10–Dec. 8, the supernova remained at R = 17.1 ± 0.1. A foreground star (R ∼ 16.0) superposed on the bulge of the host galaxy has position end figures 43^{s}.70, 03′′.2.

SUPERNOVAE CCD magnitudes (unfiltered unless otherwise noted) by L. Kiss, K.

Sárneczky, and K. Sziládi, at Piszkéstető, Hungary: SN 1999el in NGC 6951, 1999 Dec. 31.786 UT, 16.4 ± 0.1. SN 1999em in NGC 1637, Dec. 31.870, 14.6 ± 0.2. SN 1999ev in NGC 4274, Dec. 31.207, R = 18.2 ± 0.4. SN 1999ew in NGC 3677, Dec. 31.178, R = 17.7 ± 0.3. SN 1999fz in UGC 8164, Dec. 31.200, R = 17.9 ± 0.3. SN 1999gb in NGC 2532, Dec. 31.213, R = 16.8 ± 0.2. SN 1999gd in NGC 2623, Dec. 31.219, R = 17.5 ± 0.3. SN 1999gi in NGC 3184, Dec. 31.172, R = 14.6 ± 0.1. SN 1999gl in NGC 317B, Dec. 31.792, 17.8 ± 0.3. SN 1999go in NGC 1376, Dec. 31.886, 16.6 ± 0.1. SN 1999gp in UGC 1993, Dec. 29.693, R = 16.7 ± 0.2.

2000 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green